jinqx:

Triangle GHI is similar to triangle JKL. If angle H measures 60 degrees and angle L measures 25 degrees, what is the measure of angle J? (1 point) Question 14 options: 1) 25 2) 35 3) 60 4) 95

Imagine:

Any Diagrams?

jinqx:

Nope, that's all they gave me for this one

Imagine:

Okay.

Imagine:

So, It has to add to 180, Correct?

jinqx:

I believe so since it's a triangle

Imagine:

Due to it being a triangle, correct.

Imagine:

So, The math equation would look like 180-85=? That will help you solve and get your answer.

jinqx:

oh I think I see what you mean, give me just a sec

Imagine:

Due to 60+25=85

Imagine:

180-85=?= Angle "J"

jinqx:

if the triangles are similar, does that mean they both share the same angles?

Imagine:

Correct.

jinqx:

ah, so J would measure be 95? sorry my brain is so bad today I've taken like 3 exams in the past hour

Imagine:

It's okay, That is correct. 95= Angle "J"

jinqx:

okay, thank you for your help!

Imagine:

No Problem!

Imagine:

Thank you!

jinqx:

of course! <3